OpenStudy (ingah):

one model of earth population growth is p(t) equals 64/1 plus 11e^-.08t Will medal

OpenStudy (jdoe0001):

oooooooooook. noted

OpenStudy (john_es):

Any question?

OpenStudy (jdoe0001):

another model of earht is Heidi Klum btw now, she doesn't have any exponents, but she does decently in her job

OpenStudy (jdoe0001):

earth even

OpenStudy (john_es):

XD

OpenStudy (ingah):

@John_ES One model of earth population growth is p(t)=64/(1+11e^-.08t) where t is measured in the years since 1990, and p is measured in billions of people. which of the following statements are true? A:The population of Earth will grow exponentially for a while but then start to decrease B: In 1991 there were 5.74 people according to this model C: The carrying capacity of Earth is 64 billion people D: The population of Earth is growing at a rate of just under 8% per year

OpenStudy (john_es):

Well, watching the function you see that \[t\rightarrow \infty \Rightarrow p(t)\rightarrow 64\] Ok?

OpenStudy (john_es):

So the first question cannot be true. There is no exponential decay, just the model says the population will reach a constant value. Ok?

OpenStudy (ingah):

Okie

OpenStudy (john_es):

For the second, as t is measured in years taking the zero at 1990, you must evaluete the function for t=1, \[p(1)=\frac{64}{1+11e^{-0.08\cdot1}}\]

OpenStudy (john_es):

What do you obtain?

OpenStudy (ingah):

74.15

OpenStudy (ingah):

So, not true then?

OpenStudy (john_es):

Not exactly, the value is 5.7377 that is 5.74, so it is true.

OpenStudy (ingah):

Oh, I didn't put parentheses around the bottom one.

OpenStudy (john_es):

For the C question, we know from the A question that the there is a constant limit value whose value is 64. So, it is true too.

OpenStudy (ingah):

Okie

OpenStudy (john_es):

Now, the last question. They ask for the growing rate, so we must calculate the rate. To do it you must do the following: \[p(2)-p(1)\] This gives you the growth in one year. They gives a number in %, so you must compare with the original and multipliy by 100: \[\frac{p(2)-p(1)}{p(1)}\cdot1000\] Try to calculate and tell me your result.

OpenStudy (john_es):

Sorry, it is not a 1000 is a 100

OpenStudy (ingah):

I'm really confused

OpenStudy (john_es):

Ok, for the growth is just calculate the growth with the formula they provide, \[p(1)=5.74\] the previous result, And \[p(2)=6.17\]Ok?

OpenStudy (ingah):

So it would be just under 8%

OpenStudy (john_es):

So the population increase in \[p(2)-p(1)=0.43\] Now you want to calculate which percentage represents this number, so \[\frac{0.43}{p(1)}\cdot 100=7.52\approx 8\%\] So it is true.

OpenStudy (ingah):

So, B: In 1991 there were 5.74 people according to this model, C: The carrying capacity of Earth is 64 billion people & D: The population of Earth is growing at a rate of just under 8% per year are all correct, right?

OpenStudy (john_es):

Correct!

OpenStudy (ingah):

Thank you!